Suppose that the following schedule summarizes the sales (demand) and cost situation confronting a Sweezy oligopoly firm:

Price                                        $8       $10       $12       $14       $16       $17       $18       $19            $20

Quantity Demanded                       9          8          7          6          5          4          3          2                1

Total Cost                                 $128   $110      $94       $80       $68       $58       $50       $44            $40

a. Graph the demand, marginal revenue, and marginal cost curves facing the firm in a single graph.

b. Identify as q* the profit-maximizing rate of output on the same graph as in a. above.

c. What is the profit-maximizing rate of output given the information in the table above?

d. What is maximum profit for this firm?
Answer #1

A).

consider the given problem here we have given the data on demand curve and the TC. So, here “TR” is the product of “P” and “Q”, => we can get the TR for all “Q” and the MR is the additional revenue from the production of additional “q”, => MR(n) = TR(n)-TR(n-1). Now, the MC is the additional cost from the additional production, => MC(n) = TC(n) – TC(n-1). Consider the following fig. show the demand curve, MR and MC curve.

B, C).

So, here the optimum “q” is “5 units”. Since for “q=5” the “MR<MC”, => the production of additional unit is profitable but for “q=6” the “MC > MR”, => the producer should reduce production, => the optimum production is “5 units”.

D).

So, here the profit is given by, => A = TR – TC, => 16*5 – 68 = $12. So, here the profit is given by “$12”.